### v2

@@ Abstract
We develop a resource-efficient methodology for measuring economic outlook in news text that combines document embeddings with synthetic training data generated by large language models. Applied to 27 million news articles, the resulting indicator significantly improves GDP growth forecast accuracy and captures sentiment shifts weeks before official releases, proving particularly valuable during crises. The indicator outperforms both survey-based benchmarks and traditional dictionary methods and is interpretable, allowing identification of specific drivers of economic sentiment. Our approach addresses key institutional constraints: it performs sentiment classification locally, enabling analysis of proprietary news content without transmission to external services while requiring minimal computational resources compared to direct LLM classification. The methodology generalizes to other countries and restricted data environments.
@@ 1 Introduction
News articles represent a particularly valuable source of alternative data. The underlying intuition is that media coverage reflects real-time shifts in economic outlook and sentiment, often capturing developments well before they appear in official statistics or survey-based indicators (e.g., tetlock:2007; bybee:2024). A sudden policy announcement or market shock, for instance, generates immediate news coverage but may take weeks to surface in consumer confidence or business sentiment surveys. Recent advances in natural language processing have made it feasible to extract timely economic signals from news streams at scale, with studies demonstrating that such signals can meaningfully improve the precision of macroeconomic forecasts compared to traditional indicators alone (e.g., hirschenbuehl:2021; woloszko:2024; kalamara:2022; ashwin:2024).
Large language models (LLMs) have emerged as a powerful tool for extracting information from text (e.g., ash:2025; kwon:2024; kwon:2025). Unlike traditional keyword-matching approaches, LLMs capture complex semantic structures and contextual nuances in language. However, their application faces significant practical constraints. State-of-the-art models require substantial computational resources, whether through internal hosting infrastructure or external services–a challenge for many institutions. Moreover, data usage agreements (DUAs) from news providers often restrict transmitting data to external services, while internal infrastructure typically cannot support the most advanced models.
@@
Our approach uses a multi-step process that strategically combines different models to balance performance and efficiency. First, news articles are converted into high-dimensional document embeddings using a specialized, computationally efficient transformer model. Second, we use a state-of-the-art LLM to generate synthetic articles that mimic typical news articles content while conveying clearly positive or negative economic outlook. These synthetic articles are embedded and used to train a logistic regression model, which is then applied to real news article embeddings to produce economic outlook scores. The final NEOS indicator aggregates these scores across all articles in a given month. This architecture is highly resource-efficient: we estimate that operating costs of our approach are several orders of magnitude lower compared to directly assessing economic outlook using a commercial LLM, while maintaining the semantic understanding capabilities of advanced language models.
@@
Our methodology also addresses several limitations in existing research. Previous studies often analyze sentiment from single–such as the Wall Street Journal–or only a few outlets potentially missing important dimensions of public sentiment (e.g., tetlock:2007; garcia:2013). We demonstrate meaningful added value from incorporating diverse publication sources: expert publications and broader media outlets capture different facets of sentiment, and combining both provides a more balanced assessment of economic outlook.
@@
We evaluate NEOS through a pseudo-out-of-sample forecasting study of GDP growth. Our results show that NEOS significantly outperforms both survey-based benchmark indicators and traditional dictionary-based approaches. Timelier versions of NEOS, computed from partial monthly data, capture sentiment shifts weeks before official releases, proving particularly valuable during times of crises. Unlike dictionary methods that rely on simple keyword matching, NEOS captures the semantic meaning of articles, enabling deeper understanding of economic outlook. These findings align with recent research emphasizing the limitations of dictionary-based sentiment classification: while computationally efficient, such methods lack accuracy compared to transformer models and large language models (e.g., audrino:2024; kirtac:2024; klahn:2025). [^1] Previous studies by binsbergen:2024 and barbaglia:2025 demonstrate the value of integrating news-based indicators with machine learning techniques to improve forecast accuracy.
Our work relates closely to recent studies leveraging LLMs for economic sentiment analysis and forecasting, though with important distinctions. kwon:2025 extract growth and inflation sentiment indices by directly parsing press narratives using GPT-4 and GPT-5, decomposing aggregate sentiment into demand, supply, and structural drivers. While their approach improves forecasting performance, our method offers greater efficiency through embeddings that can run on on-premises systems without requiring large cloud-based infrastructure. buckmann:2025 similarly use logistic regression on embeddings from small generative models for sentiment analysis, highlighting advantages in privacy, availability, and cost. Our study extends this work through a two-sample approach using synthetic documents alongside real articles, benefiting from state-of-the-art LLMs while avoiding manual labeling. Finally, seiler:2025 computes an economic sentiment indicator from Swiss business tendency surveys using a BERT-based model. While that indicator tracks the business cycle well, our method leverages a richer, more diverse article corpus and incorporates synthetic documents to enhance robustness.
The remainder of this paper proceeds as follows. Section 2 introduces our data. Section 3 outlines the methodology for computing NEOS and extracting insights into its drivers. Section 4 evaluates NEOS’s forecasting performance and illustrates key drivers of the indicator. Section 5 presents sensitivity analyses validating our methodological design choices. Section 6 concludes.
@@ 2 Data
The Swissdox database covers almost all publications by Swiss media outlets in recent years, with the exception of Italian-language Swiss media. Therefore, our analysis focuses on articles published in German and French. We also discard some very specific media outlets that presumably do not contain any relevant articles (e.g. TV guides and various magazines). In addition, we only concentrate on print and online news articles, excluding radio and TV transcripts. Our analysis spans from January 1999 until November 2025. Thus, we use all news articles, print and online, in German and French, from 159 media outlets, amounting to roughly 27 million articles. The database is growing at a rate of about 100,000 articles per month, with new content typically becoming available two business days after publication.
We restrict our analysis to articles published from January 1999 onward to avoid the effects of compositional changes in the early years of the archive. The raw text contains markup to identify elements of an article (e.g. title, body, paragraphs, tables, boxes). During pre-processing, we remove all tables and boxes, remove all markup tags and concatenate the title and body of the article.
@@ 3 Methodology
This section describes the individual steps involved in computing NEOS. Figure 2 gives an overview. We use an efficient and scalable two-sample estimation approach to generate the indicator. As illustrated in Figure 2, we leverage both the Swissdox news articles (Section 2, top stream in Figure 2), and synthetic articles generated by an LLM (Section 3.3, bottom stream in Figure 2). We run both sets of articles through a small encoder-only embedding model to get numeric representations (Section 3.1). Using these embeddings, we select articles from the Swissdox data that relate to the state of the economy (see Section 3.2). We use only the synthetic data sample to train the sentiment model (Section 3.4). We then take the fitted model to predict the sentiment scores of the relevant articles in the Swissdox data and aggregate these scores to compute NEOS (Section 3.5). In addition, we decompose NEOS with different methods to ensure interpretability (Section 3.6). We detail each step below.
@@ 3.1 Generate document embeddings
To transform article texts into a numerical format suitable for analysis, we use the jina-embeddings-v3 model developed by JinaAI (sturua:2025). This encoder-only embedding model converts each article into a 1024-dimensional vector that captures the semantic meaning of the content, preserving the overall message and tone of the article. The model is multilingual, i.e., capable of processing both German and French articles, and can handle relatively long documents with up to 8192 input tokens.
Generating embeddings has several advantages compared to directly using LLMs to classify the news articles. Although large decoder-only models (GPT, Llama, etc.) can be used to classify sentiment in texts directly, these models typically contain tens or hundreds of billions of parameters, making them computationally expensive and slow for inference, whereas dedicated embedding models like jina-embeddings-v3 are significantly more compact and efficient. The embeddings preserve the semantic meaning, which means that the overall message and tone of a document is captured in the embedding, and can be used to build a dedicated sentiment model.
We estimate that computing an economic outlook indicator similar to NEOS using ChatGPT-4o to directly analyze the economic outlook of all news articles would have cost over \$30,000 at the time NEOS was developed. In contrast, NEOS operates at a much lower cost. It can be run on consumer-grade hardware, and when deployed on an external platform such as AWS Cloud, the computation of NEOS can be completed for around \$100, based on the costs at the time of its development. [^2] For the computation of NEOS, we used a single internally hosted NVIDIA L40S GPU with 48GB of memory.
In addition, large generative models, such as ChatGPT-4o, are commercial offerings which require data to be transmitted to external API endpoints. The licensing terms offered by the providers are typically irreconcilable with common data use agreements (DUAs) or institutional data protection requirements. This effectively prohibits processing sensitive or proprietary data. The news articles we use are subject to a restrictive DUA, so on-premise deployment is essential to maintain data sovereignty and avoid DUA violations. Licensing terms offered by other news data providers (Bloomberg, Reuters, Dow Jones Factiva) are similarly restrictive. The smaller footprint of jina-embeddings-v3 makes it feasible to deploy it on modest hardware infrastructure, while large generative models would require substantial GPU resources that are infeasible for most on-premise environments.
@@ 3.2 Select relevant articles
In order to isolate the articles relevant to the state of the economy, we employ a classification approach. Some media outlets within the Swissdox@LiRI database provide information about the section in which an article was published. We assume that articles in sections like “Business,” “Markets,” and “Economics” are likely to contain information about the state of the economy. Using the embeddings of these articles, we train a neural network model to classify whether a given article is related to economics.
This model is then applied to the entire database to filter irrelevant articles, leaving us with a refined set of articles covering economics. In total, we retain 2.7 million relevant articles in German and 0.4 million in French. The number of relevant articles per month increased over time, from around 3,000 articles in January 1999 to a peak of approximately 17,000 articles per month during the COVID-19 pandemic, and to around 13,000 articles in November 2025. Figure 3 shows the distribution of these articles over time, segmented by publication type. By selecting only articles relevant to the state of the economy, we improve the signal in the data.
To assess the performance of our neural network classifier, we randomly selected 300,000 articles, with an equal number of relevant and non-relevant articles, for training. We then tested the classifier on an additional set of 150,000 articles, which were not used during training and also equally split between relevant and non-relevant articles, to evaluate performance on a hold-out test set. We settle on a classification threshold of 80%, trading off some accuracy for gains in precision. The final model achieves 93% accuracy and 95% precision on the test data, demonstrating the strong performance of our classifier. We then use the trained classifier to label and select all relevant articles in the full data.
@@ 3.3 Generate synthetic articles
To avoid the time-intensive, error-prone and subjective process of reading and manually labeling the sentiment of parts or even all relevant articles, we leverage a state-of-the art LLM to generate synthetic articles conveying clearly positive or negative economic outlooks. The synthetic articles are created using Claude 3.5 Sonnet (Anthropic2024). We generate in total 256 synthetic articles, 128 articles that convey a stereotypical positive economic outlook and 128 articles that convey a stereotypical negative economic outlook. We instruct the LLM to generate the synthetic articles in such a way that they mimic a typical news article in style and structure. We specify that some synthetic articles should cover the economy in general, while others should cover specific parts of the economy, i.e. financial markets, the labor market, real estate market, international trade, consumption, business situation and macroeconomic outlook. An example prompt, along with a positive and negative synthetic article, is provided in Appendix A.1. We assign a score of one to the synthetic articles with a positive outlook and a score of zero to the synthetic articles with a negative outlook.
The synthetic articles are then converted into document embeddings using the jina-embeddings-v3 model described in Section 3.1. For exploratory analysis and visualization, we apply UMAP (mcinnes:2020), a dimensionality reduction approach, to reduce the high-dimensional embeddings to two dimensions. Figure 4 plots the two-dimensional representation of the 256 synthetic articles. We color the articles with positive outlook green and the ones with negative outlook red. We observe two distinct clusters corresponding to positive and negative outlook. This separation confirms that the embeddings effectively capture sentiment.
Using synthetic articles has several advantages over manual labeling of real articles. Human labeling can be error-prone, difficult to reproduce, and time-intensive. When done correctly, it is also costly, as it requires multiple individuals to independently label a large number of articles in order to validate the alignment of the labels. In contrast, generating synthetic articles is cheap and produces higher quality training data, since real world articles are rarely fully positive or negative, but typically contain ambiguous wording. The use of synthetic articles allows us to clearly assign one or zero as labels to the articles and no arbitrary values in between are required.
@@
3.4 Train sentiment model
The clear separation of the synthetic articles into two distinct clusters in Figure 4 suggests that a linear classification model can be used to distinguish between positive and negative economic outlook. To achieve this, we fit a logistic regression model to the embedded synthetic articles. Given that the dimensionality of the embeddings (1024 features) exceeds the number of observations (256 articles), we apply regularized logistic regression (cessie:1992) to prevent overfitting.
3.5 Compute NEOS

### v3

@@ Abstract
We develop a resource-efficient methodology for measuring economic outlook in news text that combines document embeddings with synthetic training data generated by large language models. Applied to 27 million news articles, the resulting indicator significantly improves GDP growth forecast accuracy and captures sentiment shifts weeks before official releases, proving particularly valuable during crises. The indicator outperforms both survey-based benchmarks and traditional dictionary methods and is interpretable, allowing identification of specific drivers of economic sentiment. Our approach addresses key institutional constraints: it performs sentiment classification locally, enabling analyses of proprietary news content without transmission to external services while requiring minimal computational resources compared to direct large language model classification.
@@ 1 Introduction
News articles represent a particularly valuable source of alternative data. The underlying intuition is that media coverage reflects real-time shifts in economic outlook and sentiment, often capturing developments well before they appear in official statistics or survey-based indicators (e.g., Tetlock, 2007; Bybee et al., 2024). A sudden policy announcement or market shock, for instance, generates immediate news coverage but may take weeks to surface in consumer confidence or business sentiment surveys. Recent advances in natural language processing have made it feasible to extract timely economic signals from news streams at scale, with studies demonstrating that such signals can meaningfully improve the precision of macroeconomic forecasts compared to traditional indicators alone (e.g., Hirschbühl et al., 2021; Woloszko, 2024; Kalamara et al., 2022; Ashwin et al., 2024).
Large language models (LLMs) have emerged as a powerful tool for extracting information from text (e.g., Ash et al., 2025; Kwon et al., 2024, 2025). Unlike traditional keyword-matching approaches, LLMs capture complex semantic structures and contextual nuances in language. However, their application faces significant practical constraints. State-of-the-art models require substantial computational resources, whether through internal hosting infrastructure or external services–a challenge for many institutions. Moreover, data usage agreements (DUAs) from news providers often restrict transmitting data to external services, while internal infrastructure typically cannot support the most advanced models.
@@
Our approach uses a multi-step process that strategically combines different models to balance performance and efficiency. First, news articles are converted into high-dimensional document embeddings using a specialized, computationally efficient transformer model. Second, we use a state-of-the-art LLM to generate synthetic articles that mimic typical news article content while conveying a clearly positive or negative economic outlook. These synthetic articles are embedded and used to train a logistic regression model, which is then applied to real news article embeddings to produce economic outlook scores. The final NEOS indicator aggregates these scores across all articles in a given month. This architecture is highly resource-efficient: we estimate that the operating costs of our approach are several orders of magnitude lower compared to directly assessing economic outlook based on the same corpus of news articles using a commercial LLM, while maintaining the semantic understanding capabilities of advanced language models.
@@
Our methodology also addresses several limitations of the existing research. Previous studies have often analyzed sentiment from single outlets–such as the Wall Street Journal–or only a few outlets, potentially missing important dimensions of public sentiment (e.g., Tetlock, 2007; García, 2013). We demonstrate meaningful added value from incorporating diverse publication sources: expert publications and broader media outlets capture different facets of sentiment, and combining both provides a more balanced assessment of economic outlook.
@@
We evaluate NEOS through a pseudo-out-of-sample forecasting study of GDP growth. Our results show that NEOS significantly outperforms both survey-based benchmark indicators and traditional dictionary-based approaches. Timelier versions of NEOS, computed from partial monthly data, capture sentiment shifts weeks before official releases, proving particularly valuable during times of crises. Unlike dictionary methods that rely on simple keyword matching, NEOS captures the semantic meaning of articles, enabling deeper understanding of economic outlook. These findings align with recent research emphasizing the limitations of dictionary-based sentiment classification: while computationally efficient, such methods lack accuracy compared to transformer models and large language models (e.g., Audrino et al., 2024; Kirtac and Germano, 2024; Klähn et al., 2025). [^1] Previous studies by van Binsbergen et al. (2024) and Barbaglia et al. (2025) demonstrated the value of integrating news-based indicators with machine learning techniques to improve forecast accuracy.
Our work relates closely to recent studies leveraging LLMs for economic sentiment analysis and forecasting, though with important distinctions. Kwon et al. (2025) extracted growth and inflation sentiment indices by directly parsing press narratives using GPT-4 and GPT-5, decomposing aggregate sentiment into demand, supply, and structural drivers. While their approach improves forecasting performance, our method offers greater efficiency through embeddings that can run on on-premises systems without requiring large cloud-based infrastructure. Buckmann and Hill (2025) similarly used logistic regression on embeddings from small generative models for sentiment analysis, highlighting advantages in privacy, availability, and cost. Our study extends this work through a two-sample approach using synthetic documents alongside real articles, benefiting from state-of-the-art LLMs while avoiding manual labeling. Finally, Seiler (2025) computed an economic sentiment indicator from Swiss business tendency surveys using a BERT-based model. While that indicator tracks the business cycle well, our method leverages a richer, more diverse article corpus and incorporates synthetic documents to enable comparability.
The remainder of this paper proceeds as follows. Section 2 introduces our data. Section 3 outlines the methodology for computing NEOS and extracting insights into its drivers. Section 4 evaluates NEOS’s forecasting performance and illustrates key drivers of the indicator. Section 5 presents sensitivity analyses validating our methodological design choices. Section 6 concludes the study.
@@ 2 Data
The Swissdox database covers almost all publications by Swiss media outlets in recent years, with the exception of Italian-language Swiss media. Therefore, our analysis focuses on articles published in German and French. We also discard some very specific media outlets that presumably do not contain any relevant articles (e.g., TV guides and various magazines). In addition, we only concentrate on print and online news articles, excluding radio and TV transcripts. Our analysis spans from January 1999 until November 2025. Thus, we use all news articles, print and online, in German and French, from 159 media outlets, amounting to roughly 27 million articles. The database is growing at a rate of approximately 100,000 articles per month, with new content typically becoming available two business days after publication.
We restrict our analysis to articles published from January 1999 onward to avoid the effects of compositional changes in the early years of the archive. The raw text contains markup to identify elements of an article (e.g., title, body, paragraphs, tables, boxes). During preprocessing, we remove all tables and boxes, remove all markup tags and concatenate the title and body of the article.
@@ 3 Methodology
This section describes the individual steps involved in computing NEOS. Figure 2 gives an overview. We use an efficient and scalable two-sample estimation approach to generate the indicator. As illustrated in Figure 2, we leverage both the Swissdox news articles (Section 2, i.e., the top stream in Figure 2), and synthetic articles generated by an LLM (Section 3.3, i.e., the bottom stream in Figure 2). We run both sets of articles through a small encoder-only embedding model to obtain the numeric representations (Section 3.1). Using these embeddings, we select articles from the Swissdox data that relate to the state of the economy (see Section 3.2). We use only the synthetic data sample to train the sentiment model (Section 3.4). We then take the fitted model to predict the sentiment scores of the relevant articles in the Swissdox data and aggregate these scores to compute NEOS (Section 3.5). In addition, we decompose NEOS with different methods to ensure interpretability (Section 3.6). We detail each step below.
@@ 3.1 Generate document embeddings
To transform article texts into a numerical format suitable for analysis, we use the jina-embeddings-v3 model developed by JinaAI (Sturua et al., 2025). This encoder-only embedding model converts each article into a 1024-dimensional vector that captures the semantic meaning of the content, preserving the overall message and tone of the article. We choose jina-embeddings-v3 because the model is multilingual, i.e., capable of processing both German and French articles, and can handle relatively long documents with up to 8192 input tokens. Previous generation embedding models like BERT only support short input sequences of 512 tokens, which are insufficient to generate high quality embeddings of longer content like news articles. At the same time, jina-embeddings-v3 is still a comparatively small model with only 570M parameters, allowing for fast inference on moderate hardware.
Generating embeddings has several advantages compared to directly using LLMs to classify the news articles. Although large decoder-only models (GPT, Llama, etc.) can be used to classify sentiment in texts directly, these models typically contain tens or hundreds of billions of parameters, making them computationally expensive and slow for inference, whereas dedicated embedding models like jina-embeddings-v3 are significantly more compact and efficient. The embeddings preserve the semantic meaning, which means that the overall message and tone of a document is captured in the embeddings and can be used to build a dedicated sentiment model.
We estimate that computing an economic outlook indicator similar to NEOS using ChatGPT-4o to directly analyze the economic outlook of all news articles would have cost over \$30,000 at the time that NEOS was developed. In contrast, NEOS operates at a much lower cost. It can be run on consumer-grade hardware, and if deployment on an external platform such as AWS Cloud were allowed, the computation of NEOS could be completed for approximately \$100, based on the costs at the time of development. [^2] For development, we used a single local NVIDIA L40S GPU with 48GB of memory. Note that the Swissdox@LiRI data are subject to restrictive DUAs and cannot be analyzed on external platforms.
In addition, large generative models, such as ChatGPT-4o, are commercial offerings which require data to be transmitted to external API endpoints. The licensing terms offered by the providers are typically irreconcilable with common data use agreements (DUAs) or institutional data protection requirements. This effectively prohibits processing sensitive or proprietary data. The news articles we use are subject to a restrictive DUA, so on-premises deployment is essential to maintain data sovereignty and avoid DUA violations. Licensing terms offered by other news data providers (Bloomberg, Reuters, Dow Jones Factiva) are similarly restrictive. The smaller footprint of jina-embeddings-v3 makes it feasible to deploy it on modest hardware infrastructure, while large generative models would require substantial GPU resources that are infeasible for most on-premises environments.
@@ 3.2 Select relevant articles
To isolate the articles relevant to the state of the economy, we employ a classification approach. Some media outlets within the Swissdox@LiRI database provide information about the section in which an article was published. We assume that articles in economics-related sections like “business”, “markets”, and “economics” are likely to contain information about the state of the economy. We consider these economics-related articles as relevant, and articles published in other sections (e.g. “sports”) non-relevant. We use the embeddings of the articles with known sections to train a neural network model to classify whether any given article is related to economics (i.e., relevant) or not.
We evaluate different candidate models to classify the relevant (economics-related) articles before settling on the neural network. To train the models, we use a balanced sample of 300,000 articles, randomly selecting 150,000 articles each from both the relevant and non-relevant articles. We then evaluate the performance of the classifiers on a test set, an additional set of 150,000 articles, which were not used during training, also equally split between relevant and non-relevant articles. In our model evaluation, we achieved approximately 92% accuracy using a linear model (L2-penalized logistic regression). A random forest also achieved 92% accuracy. We settle on a simple feedforward neural network with four hidden layers in a funnel structure (512, 256, 128 and 64 neurons), ReLU activation, L2 regularization, dropout, batch normalization and sigmoid output layer; achieving 95% accuracy.
Next, we evaluate the precision-recall tradeoff. Figure 15 in the Appendix A.6 displays the overall accuracy vs. precision for different candidate models, illustrating our choice of model and classification threshold. To improve the model’s ability to correctly identify relevant articles (precision) and limit the number of false positives, we increase the classification threshold from 0.5 to 0.8, trading off some accuracy for gains in precision. The final model achieves 93% accuracy and 95% precision on the test data, demonstrating the strong performance of our classifier. We then use the trained classifier to label and select all relevant articles in the full data.
Based on the model prediction, we select only relevant articles from the full sample, leaving us with a refined set of economics-related articles. In total, we retain 2.7 million relevant articles in German and 0.4 million relevant articles in French. The number of relevant articles per month increased over time, from approximately 3,000 articles in January 1999 to a peak of approximately 17,000 articles per month during the COVID-19 pandemic, and to approximately 13,000 articles in November 2025. Figure 3 shows the distribution of these articles over time, segmented by publication type. By selecting only articles relevant to the state of the economy, we improve the signal in the data.
@@ 3.3 Generate synthetic articles
To avoid the time-intensive, error-prone and subjective process of reading and manually labeling the sentiment of parts or even all relevant articles, we leverage a state-of-the-art LLM to generate synthetic articles conveying clearly positive or negative economic outlooks. The synthetic articles are created using Claude 3.5 Sonnet (Anthropic, 2024). We generate in total 256 synthetic articles, 128 articles of which convey a stereotypical positive economic outlook and 128 articles of which convey a stereotypical negative economic outlook. We instruct the LLM to generate the synthetic articles in such a way that they mimic a typical news article in style and structure. We specify that some synthetic articles should cover the economy in general, while others should cover specific parts of the economy, i.e., financial markets, the labor market, real estate market, international trade, consumption, business situation and macroeconomic outlook. An example prompt, along with a positive and negative synthetic article, is provided in Appendix A.1. We assign a score of one to the synthetic articles with a positive outlook and a score of zero to the synthetic articles with a negative outlook.
The synthetic articles are then converted into document embeddings using the jina-embeddings-v3 model described in Section 3.1. For exploratory analysis and visualization, we apply UMAP (McInnes et al., 2020), a dimensionality reduction approach, to reduce the high-dimensional embeddings to two dimensions. Figure 4 plots the two-dimensional representation of the 256 synthetic articles. We color the articles with positive outlook green and the ones with negative outlook red. We observe two distinct clusters corresponding to positive and negative outlook. This separation confirms that the embeddings effectively capture sentiment.
Using synthetic articles has several advantages over manual labeling of real articles. Human labeling can be error-prone, difficult to reproduce, and time-intensive. When performed correctly, it is also costly, as it requires multiple individuals to independently label a large number of articles to validate the alignment of the labels. In contrast, generating synthetic articles is cheap and produces higher quality training data, since real world articles are rarely fully positive or negative, but typically contain ambiguous wording. The use of synthetic articles allows us to clearly assign one or zero as labels to the articles and no arbitrary values in between are required.
@@
3.4 Train the sentiment model
The clear separation of the synthetic articles into two distinct clusters in Figure 4 suggests that a linear classification model can be used to distinguish between positive and negative economic outlook. To achieve this, we fit a logistic regression model to the embedded synthetic articles. Given that the dimensionality of the embeddings (1024 features) exceeds the number of observations (256 articles), we apply regularized logistic regression (Cessie and Houwelingen, 1992) to prevent overfitting.
3.5 Measure economic outlook